OpenStudy (anonymous):

Newton method: Volume of tank -> V = (4/3)(pi)(r^3) + (pi)(r^2)(h) V = 50, h = 4, so 50 = 4/3(pi)(r^3) + 4(pi)r^2 or 50/(pi) = 4/3(r^3) + 4(r^2) Take r1 = 1.5 as initial guess, find f(r) and f'(r)

OpenStudy (bahrom7893):

@Zarkon can u help?

OpenStudy (bahrom7893):

okay i'll try myself.

OpenStudy (anonymous):

The f(r), is that just the equation in standard form, or should I also sub in the value 1.5 to get an actual number?

OpenStudy (bahrom7893):

no not yet: x_(n+1) = x_n - f(x_n)/f'(x_n)

OpenStudy (bahrom7893):

f(r) = (4/3)(pi)(1.5^3) + (pi)(1.5^2)(4)

OpenStudy (bahrom7893):

f'(r) = [ (4/3)(pi)(r^3) + 4(pi)(r^2) ] ' = 4pi*r^2 + 8pi*r

OpenStudy (bahrom7893):

f ' (r) = 4pi*(1.5)^2 + 8pi*(1.5)

OpenStudy (anonymous):

So am I just writing out the equation for f(r) and subbing in 1.5, and then for f'(r) I find the derivative of the first equation and THEN sub in 1.5?

OpenStudy (bahrom7893):

yes

OpenStudy (bahrom7893):

but first sub in h for both

OpenStudy (bahrom7893):

do u need more help?

OpenStudy (anonymous):

Umm, give me a moment to look it over I got distracted by something.

OpenStudy (anonymous):

I don't understand this: x_(n+1) = x_n - f(x_n)/f'(x_n) Where is the n variable coming from?

OpenStudy (bahrom7893):

it's x sub n

OpenStudy (bahrom7893):

x sub n is the first guess which is your r x sub n+1 is the next guess.

OpenStudy (bahrom7893):